OpenStudy (anonymous):

Solve the equation. 78 = -2(m + 3) + m

OpenStudy (anonymous):

1. Expand it first 78 = -2m-6+9 2. Simplify −2m−6+m to −m−6 78=-m-6 3.Add 6 to both sides 78+6=-m 4.Simplify 78+6 to 84 84=-m 5.Multiply both sides by 1 -84=m 6.Switch sides m+-84

OpenStudy (anonymous):

final answer is m=-84

OpenStudy (anonymous):